# LLM Pruning and Distillation in Practice: The Minitron Approach

### v1

@@ 1 Introduction
Recent work [1] has demonstrated the effectiveness of combining weight pruning with knowledge distillation to significantly reduce the cost of training LLM model families. Here, only the biggest model in the family is trained from scratch; other models are obtained by successively pruning the bigger model(s) and then performing knowledge distillation to recover the accuracy of pruned models.
@@ 2.2 Retraining with Distillation
We use the term retraining to refer to the accuracy recovery process following pruning. In this work, we explore two retraining strategies: (1) conventional training, leveraging ground truth labels, and (2) knowledge distillation using supervision from the unpruned model (teacher). Knowledge Distillation (KD) involves transfer of knowledge from a larger or more complex model called the teacher to a smaller/simpler model called the student. The knowledge transfer is achieved by having the student model mimic the output and/or the intermediate states of the teacher model. In our case, the uncompressed and pruned models correspond to the teacher and student, respectively. For distillation, we follow best practices from our previous work [1] and use forward KL Divergence loss [7] on the teacher and student logits only. This is illustrated in Figure 3.
@@ 3.1 Pre-training
We use the Nemotron-4 curated continued training dataset(CT) [8]  [9] for all our experiments.
@@ 6.0.2 Mistral NeMo 12B to MN-Minitron-8B:
    Depth pruning boosts throughput, achieving $\sim 2.7\times$∼ 2.7 × speedup over Llama-3.1 8B, while width pruning provides $\sim 1.7\times$∼ 1.7 × speed up.

### v2

@@ 1 Introduction
Recent work [1] has demonstrated the effectiveness of combining weight pruning with knowledge distillation to significantly reduce the cost of training LLM model families. Here, only the biggest model in the family is trained from scratch; other models are obtained by successively pruning the bigger model(s) and then performing knowledge distillation [3] to recover the accuracy of pruned models.
@@ 2.2 Retraining with Distillation
We use the term retraining to refer to the accuracy recovery process following pruning. In this work, we explore two retraining strategies: (1) conventional training, leveraging ground truth labels, and (2) knowledge distillation using supervision from the unpruned model (teacher). Knowledge Distillation (KD) [3] involves transfer of knowledge from a larger or more complex model called the teacher to a smaller/simpler model called the student. The knowledge transfer is achieved by having the student model mimic the output and/or the intermediate states of the teacher model. In our case, the uncompressed and pruned models correspond to the teacher and student, respectively. For distillation, we follow best practices from our previous work [1] and use forward KL Divergence loss [8] on the teacher and student logits only (following the [3]). This is illustrated in Figure 3.
@@ 3.1 Pre-training
We use the Nemotron-4 curated continued training (CT) dataset [9]  [10] for all our pruning and distillation experiments.
@@ 5.2 Instruct Models
5.3 Runtime Performance Analysis
We optimized the Llama 3.1 8B and Llama-3.1-Minitron 4B variants with NVIDIA TensorRT-LLM, an open-source toolkit for optimized LLM inference.
Figure 10: TensorRT-LLM FP8 throughput comparison for the Llama-3.1-Minitron-4B models with the Llama 3.1 8B model w.r.t. increasing input and output sequence lengths.
Figure 10 shows the throughput in requests per second for the various models in FP8 precision obtained on a single H100 80 GB GPU. Different use cases are represented by increasing input sequence length/output sequence length (ISL/OSL) combinations, at a batch size of 32 and 64 for the 8B-12B models and the 4B models respectively. The smaller memory footprint of the 4B model allows for larger batches. We notice that Llama-3.1-Minitron-4B-Depth is the fastest, achieving an average throughput improvement of $2.7\times$2.7 × over Llama 3.1 8B; the Llama-3.1-Minitron-4B-Width variant achieves an average throughput improvement of $1.8\times$1.8 × over Llama 3.1 8B. Compared to BF16, we notice that FP8 delivers a performance boost of $1.4\times$1.4 ×.
@@ 6.0.2 Mistral NeMo 12B to MN-Minitron-8B:
    Depth pruning boosts throughput, achieving $2.7\times$2.7 × speedup over Llama-3.1 8B, while width pruning provides $1.7\times$1.7 × speedup.